OpenStudy (anonymous):

Simplify the expression. x^2 • x^3

jimthompson5910 (jim_thompson5910):

The bases are both the same (both are x) When you multiply two exponential expressions with the same base, you add the exponents. This means x^2 * x^3 = x^(2+3) x^2 * x^3 = x^5

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

yw